# FitzHugh-Nagumo equation: from global dynamics to the slow-fast system

### v1

@@ 5. The phase portraits in the Poincaré disc
    a saddle point for
    $\{(a,b,c)\big{|}b>0,\,c<0\}\cup\{(a,b,c)\big{|}b>0,\,c>0\}.$ (5.13)
@@
    an unstable node for
    $\{(a,b,c)\big{|}b<0,\,c<0\}.$ (5.14)
@@
    a stable node for
    $\{(a,b,c)\big{|}b<0,\,c>0\}\cup\{(a,b,c)\big{|}b=0\}.$ (5.15)
The final step of the proof follows from the description of Subsection 5.1. ∎
(a) Region a
(b) Region b
(c) Region c
(d) Region d
(e) Region e
(f) Region f
(g) Region g
(h) Region h
(i) Region i
(j) Region j
(k) Region k
(l) Region l
(m) Region m
(n) Region n
(o) Region o
(p) Region p
Figure 8. The global phase portraits of (3.1) for Case A, which is associated with system (3.1) when $a=0,$ associated with regions a–j of Figure (3)(a).
(a) Region I
(b) Region II
(c) Region III
(d) Region IV
Figure 9. The global phase portraits of (3.1) for Case B, which is associated with system (3.1) when $b=0,$ corresponding to regions I–IV of Figure 4(b), in the case where the eigenvalues of (3.5) at $E_{1}$ are complex non-real.
5.1. Description of Figures 11, 12, 13, 14, 15 and 16 in Appendix 2
In order to plot the complete phase diagram in the compactified space for the Cases A, B and C in Figures 8, 9 and 10, we need to carefully go back through each step of the blow-up procedure, explaining the transformations (blow-ups, directional transformations and coordinate maps) used at each stage, and clearly stating the resulting system at every step.
Undoing the rescaling $dt_{3}=u_{6}dt_{2}$ the phase portrait depicted in Figures 11(a), 12(a), 13(a), 14(a), 15(a) and 16(a) yields the local phase portrait at the origin of (5.10), which is topologically equivalent to that of Figures 11(b), 12(b), 13(b), 14(b), 15(b) and 16(b), respectively.
Going back through the change of variables $(u_{5},v_{5})=(u_{6},u_{6}v_{6}),$ the phase portrait depicted in Figures 11(b), 12(b), 13(b), 14(b), 15(b) and 16(b) yields the local phase portrait at the origin of (5.9), which is topologically equivalent to that of Figures 11(c), 12(c), 13(c), (14)(c), 15(c) and 16(c), respectively.
Going back through the change of variables $(u_{4},v_{4})=(u_{5}-v_{5},v_{5}),$ the phase portrait depicted in Figures 11(c), 12(c), 13(c), 14(c), 15(c) and 16(c), yields the local phase portrait at the origin of (5.8), which is topologically equivalent to that of Figures 11(d), 12(d), 13(d), 14(d), 15(d) 16(d), respectively.
Going back through the change of variables $(u_{3},v_{3})=(u_{4},1+v_{4})$ the phase portrait depicted in Figures 11(d), 12(d), 13(d), 14(d), 15(d) and 16(d) yields the local phase portrait at the origin of system (5.7), which is topologically equivalent to that of Figures 11(e), 12(e), 13(e), 14(e), 15(e) and 16(e), respectively.
Undoing the rescaling $dt_{2}=u_{3}dt_{1}$ the phase portrait depicted in Figures 11(e), 12(e), 13(e), 14(e), 15(e) and 16(e) yields the local phase portrait at the origin of (5.6), which is topologically equivalent to that of Figures 11(f) and 12(f), 13(f), 14(f), 15(f) and 16(f), respectively.
Going back through the change of variables $(u_{2},v_{2})\rightarrow(u_{3},v_{3}/u_{3}),$ the phase portrait depicted in Figures 11(f), 12(f), 13(f), 14(f), 15(f) and 16(f), yields the local phase portrait at the origin of system 5.5, which is topologically equivalent to that of Figures 11(g), 12(g), 13(g), 14(g), 15(g) and 16(g), respectively.
Going back through the change of variables $(u_{1},v_{1})=(u_{2}-v_{2},v_{2}),$ the phase portrait depicted in Figures 11(g), 12(g), 13(g), 14(g), 15(g) and 16(g) yields the local phase portrait at the origin of (5.4), which is topologically equivalent to that of Figures 11(h), 12(h), 13(h), 14(h), 15(h) and 16(h), respectively.
Undoing the rescaling $dt_{1}=u_{1}dt$ the phase portrait depicted in Figures 11(h), 12(h), 13(h), 14(h), 15(h) and 16(h) yields the local phase portrait at the origin of (5.3), which is topologically equivalent to that of Figures 11(i), 12(i), 13(i), 14(i), 15(i) and 16(i), respectively.
Going back through the change of variables $(u,v)\rightarrow(u_{1},v_{1}/u_{1}),$ the phase portrait depicted in Figures 11(i), 12(i), 13(i), 14(i), 15(i) and 16(i) yields the local phase portrait at the origin of (5.2), which is topologically equivalent to that of Figures 11(j), 12(j), 13(j), 14(j), 15(j) and 16(j), respectively.
This allows us to plot the complete phase diagram in the compactified space for the Cases A, B and C in Figures 8, 9 and 10. The local phase portraits of the blow-down at the origin of the different local charts have been plotted in Figures 1116.
(a) $c<-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$
(b) $-\sqrt{\frac{b}{1-{x^{*}}^{2}}}<c<0$
(c) $0<c<\sqrt{\frac{b}{1-{x^{*}}^{2}}}$
(d) $c>\sqrt{\frac{b}{1-{x^{*}}^{2}}}$
@@ 6. Discussion and concluding remark
Although the FHN equations were created as a simplified model for nerve impulse, they have also been intensively studied for purely mathematical reasons [12, 13, 10, 23] because they provide a very simple example of equations that exhibit rich dynamics.
Results obtained on [21, pp. 180] synthesised the global bifurcation diagram for the FHN model (3.1). It has been obtained by putting together, as in a “huge puzzle”, all local bifurcation diagrams obtained in previous chapters of the latter reference. The authors have concentrated their attention to a particular parameter region relevant to physiology ($|c|>1+\sqrt{3}$).
Trying to complete and understand the bifurcation diagram of (3.1), in this paper we have discussed the finite equilibria of (3.1), as well as their bifurcations, for the following three scenarios identified in Subsection 1.4:
@@
  • Point $\textbf{Q}_{6}$: Double Homoclinic (Theorem 3.4, curve $b={DH}$);
@@
The non-hyperbolic equilibria corresponding to parameter values situated at Q (see [21]) are degenerated saddle-node (cusp) equilibria, which are attracting for $c<1$ and repelling for $c>1$. For $c=1$, following the Remark 3.5, this point might correspond to a degenerated Bogdanov-Takens of order two (with symmetry) – a kind of codimension-three bifurcation. As pointed out in [23], this would be the unique generic codimension-three local bifurcation exhibited by the FHN model (3.1) and its complete understanding is an open problem. Using the same line of argument, in Cases B and C, we have analysed the dynamics of (3.1).
@@
Based on [10, 14], in Section 4, we have studied the asymptotic case $c\to\pm\infty$, where canards are detected. An important consideration is the empirical difficulty in finding canards, since the smaller $\varepsilon$ is, the narrower the interval of $b$ (Case A) or $a$ (Case B) values for which canards appear – see Figure 7. The dynamics produced in the system around these small intervals is called the canard explosion and has been studied in [10, 14, 12].
@@
In Section 5, we have continued the analysis with the compactification of the phase portraits associated with (3.1) on the Poincaré disc. This brings additional information of the trajectories which tend to or come from infinity. We have provided phase and bifurcation diagrams in all the three cases.
Novelty versus limitations
We have considered several cases of the classical FHN system which allow the catalogue of topologically distinct phase portraits. The global dynamics seems to be determined by the local bifurcations found in Theorems 3.4, 3.8 and Proposition 3.10. Although we have not achieved complete phase portraits for all cases, we have described all topological regions worth to be analysed. Numerics suggest that, in the three scenarios under consideration, we have obtained a complete phase portrait (up to conjugacy). However, we are not claiming it analytically.
@@
We have connected the local bifurcation theory (Theorems 3.4, 3.8 and Proposition 3.10) with the asymptotic dynamics $|c|\to+\infty$ (Propositions 4.9 and 4.10). Finally, using a blow-up technique we have depicted the dynamics of (3.1) at infinity (Theorem 5.1). Our contributions do not finish the whole discussion of the bifurcation analysis of (3.1); bifurcations might make part of of high codimension phenomena.
@@
Theoretically, the maximum number of non-trivial periodic solutions of (3.1) is three [21, pp. 218]. There are regions where the dynamics is completely determined and the phase portrait is complete (Regions b, g, o and p of Figure 8 combined with Lemma 3.1) and others in which the complete analysis of the bifurcation diagram of [21, pp. 180] is still ongoing. We defer the complete analysis for a future task.
Acknowledgments
Alexandre Rodrigues has been supported by the Project CEMAPRE/REM (UIDB/05069/ 2020) financed by FCT/MCTES through national funds. Nasrin Sadri is supported by Institute for Research in Fundamental Sciences (IPM) grant 1403370046. This work is based upon research funded by Iran National Science Foundation (INSF) under Project Number 4041099. The authors are also indebted to two reviewers for the suggestions which helped to improve the readability of this manuscript.
Appendix A Blowing-down of the origin of the local chart $U_{2}$:
Figure 11. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b>0$ and $c<0.$
Figure 12. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b<0$ and $c<0.$
Figure 13. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ when $b=0,$ $c<0.$
Figure 14. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ when $b=0,$ $c>0.$
Figure 15. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b>0$ and $c>0.$
Figure 16. The local phase portraits of the blow-down at the origin of the local chart $U_{2}$ for $b<0$ and $c>0.$
@@ Appendix A Blowing-down of the origin of the local chart $U_{2}$:
Table 1. The local phase portraits corresponding to Case A, which is associated with system (3.1) when $a=0,$ for Regions 1–28 of Figure (3)(b).
@@
Table 2. The local phase portraits corresponding to Case B, associated with system (3.1) when $b=0$, for Regions 1–8 of Figure 4(a).

### v3

@@ 1. Claim 1:
      The equilibrium $D_{1}$ always is a saddle point for $X_{1}^{+}$.
@@
      The points $f^{\pm}$ are semi-hyperbolic and the dynamics of $X_{1}^{\pm}$ restricted to the center manifold $\bar{v}=\mp\frac{\sqrt{3}(b+3c^{2})\rho^{2}}{18c^{2}}$ is given by $\dot{\rho}=\frac{b}{9c}{\rho^{3}}.$
      Proof.
      A center manifold in the form of $h(\rho):=\alpha_{1}\rho^{2}+h.o.t$ must satisfy the equation
      $\dot{\rho}\frac{\partial h}{\partial\rho}-\dot{\bar{v}}\big|_{(\rho,h(\rho))}=0.$
      After shifting $f^{\pm}$ to the origin, we obtain $\alpha_{1}=\mp\frac{\sqrt{3}(b+3c^{2})}{18c^{2}}.$ Then, the dynamics of $X_{1}^{\pm}$ restricted to the center manifold $\bar{v}=\mp\frac{\sqrt{3}(b+3c^{2})\rho^{2}}{18c^{2}}$ is given by $\dot{\rho}=\frac{b}{9c}{\rho^{3}}$ (more details in [5, Theorem 2.19]). ∎
@@
      The equilibria $f^{\pm}$ are:
      • 1a.
@@
      • 2a.
        an unstable node for $\{(a,b,c)\big|b<0,\,c<0\}$ – see Figure 10(c).
      • 3a.
        a stable node for $\{(a,b,c)\big|b<0,\,c>0\}.$ – see Figure 10(d).
  2. Claim 2:
    The equilibria of system $X_{1}^{-}$ on $\rho=0$ are determined by the roots of
    $\bar{v}\left(3\bar{v}^{2}-1\right)=0.$
    There are three equilibria at $\rho=0$ denoted by $D_{2}:(0,0)$ and $g^{\pm}:\left(0,\pm\frac{\sqrt{3}}{3}\right),$ and the results of Claim 1 are valid in this case (changing $D_{1}$ and $f^{\pm}$ by $D_{2}$ and $g_{\pm}$, respectively).
    Using the transformation (5.4) with $\bar{v}=1$ and $\bar{v}=-1,$ the $y$ directional blow-up transforms system (5.3) into
    $Y_{1}^{+}:\begin{cases}\frac{d{\rho}}{d\eta}=\mathcal{O}(|\rho,\bar{u}|),\\ \\ \frac{d\bar{u}}{d\eta}=c,\end{cases}\quad\quad Y_{1}^{-}:\begin{cases}\frac{d{\rho}}{d\eta}=\mathcal{O}(|\rho,\bar{u}|),\\ \\ \frac{d\bar{u}}{d\eta}=c,\end{cases}$
    respectively, where $d\eta=\rho^{4}d\tau.$ We now need to analyze the point (0, 0) of $Y_{1}^{\pm}.$ Indeed,
  3. Claim 3:
    For the vector fields $Y_{1}^{\pm},$ point (0, 0) is a regular point (not an equilibrium).
From the Claims 1–3, we know the dynamics of the vector fields $X_{1}^{\pm}$ and $Y_{1}^{\pm}$. For the vector fields $Y_{1}^{\pm},$ $(0,0)$ is a regular point. Thus, Figure 10 illustrates the qualitative properties of the vector fields $X_{1}^{\pm}$ and $Y_{1}^{\pm}.$ By shrinking the circles in these figures to the origin (as the radii of the circles decrease to zero and the circles collapse at $(0,0)$ ), we achieve to the phase portraits shown in Figure 11. So, as shown in Figure 8, the equilibrium of chart $U_{2}$ is a saddle with four hyperbolic sectors for $b\geq 0.$ Moreover, it becomes an unstable (resp. stable) node when $b,c<0$ and $b<0,c>0,$ respectively.
(a) $b\geq 0$ and $c>0$
(b) $b\geq 0$ and $c<0$
(c) $b<0$ and $c<0$
(d) $b<0$ and $c>0$
Figure 10. Qualitative properties of vector fields $X_{1}^{\pm}$ and $Y_{1}^{\pm}.$ The axes corresponding to systems $X_{1}^{\pm}$ and $Y_{1}^{\pm}$ are $(\rho,\bar{v})$ and $(\rho,\bar{u}),$ respectively.
(a) $b\geq 0$ and $c>0$
(b) $b\geq 0$ and $c<0$
(c) $b<0$ and $c<0$
(d) $b<0$ and $c>0$
Figure 11. Local phase portraits of system (5.3) (the representation of system (3.1) in the chart $U_{2}$) at $(0,0)$.
@@ 6. Discussion and concluding remark
Although the FHN equations were originally developed as a simplified representation of nerve impulses, they have also been extensively investigated in purely mathematical purposes [12, 13, 10, 23] – they consist of a very simple model of equations with rich dynamics.
Results obtained on [21, pp. 180] synthesised the global bifurcation diagram for the FHN system (3.1). It has been obtained by putting together, as in a “huge puzzle”, all local bifurcation diagrams obtained. The authors have concentrated their attention to a particular parameter region relevant to physiology ($|c|>1+\sqrt{3}$).
Trying to complete and understand the bifurcation diagram of (3.1), in this paper we have discussed the finite equilibria of (3.1), as well as their bifurcations, for the three scenarios identified in Subsection 1.4:
@@
  • Point $\textbf{Q}_{6}$: Double Homoclinic (Theorem 3.4, curve b=DHb={DH});
@@
The non-hyperbolic equilibria associated with parameter values located at Q (see [21]) are degenerated saddle-node (cusp) equilibria, which are attracting for $c<1$ and repelling for $c>1$. For $c=1$, following the Remark 3.6, this point might correspond to a degenerated Bogdanov-Takens of order two (with symmetry) – a kind of codimension-three bifurcation. As pointed out in [23], this would be the unique generic codimension-three local bifurcation exhibited by the FHN model (3.1) and its complete understanding is an open problem. Using the same line of argument, in Cases B and C, we have analysed the dynamics of (3.1).
@@
In Section 4, we have studied the asymptotic case $c\to\pm\infty$, where canards are detected. A key point in this research is the empirical difficulty to detect canards since the smaller value of $\varepsilon$ is, the narrower the interval of $b$ (Case A) or $a$ (Case B) values for which canards appear – see Figure 7. The dynamics that emerges in the system within these narrow parameter ranges is known as the canard explosion and has been studied in [10, 14, 12].
@@
In Section 5, we have continued the analysis with the compactification of the phase portraits of (3.1) on the Poincaré disc. This brings additional information of the trajectories which come from or tend to infinity. We have provided phase and bifurcation diagrams in all the three cases.
@@
Throughout this article, we have connected the local bifurcation theory (Theorems 3.4, 3.9 and Proposition 3.10) with dynamics “near infinity” (Theorem 5.1). Our contributions do not finish the whole discussion of the bifurcation analysis of (3.1); bifurcations might make part of high codimension phenomena.
@@
Theoretically, the maximum number of non-trivial periodic solutions of (3.1) is three [21, pp. 218]. There are regions where the dynamics is completely determined and the phase portrait is complete and others in which the complete analysis of the bifurcation diagram of [21, pp. 180] is still ongoing.
7. Appendix
All phase portrait are illustrated with numerical simulations in Tables 1 and 2 using the software pplane9. Initial conditions were rigorously chosen to illustrate the characteristic behaviour in each region. Furthermore, limit cycles are indicated by the red trajectories in the phase plot. The phase portraits of Table 1 and 2 have been plotted for the parameter values,
$(c,b)=$ $(2.5,1.1),\;(2,1.2),\;(2,1.3),\;(2,1.34),\;(1.3,1.3),\;(0.5,1.1),\;(-0.5,1.1),\;(-2,1.5),$
$(-2,1.34),\;(-2,1.33),\;(-2,1.2),\;(-2.5,1.1),\;(-3,0.5),\;(-1,0.5),\;(-0.5,0.4)\;$
$(-0.1,0.9),\;(0.1,0.9),\;(0.4,0.4),\;(1.5,0.5),\;(3,0.5),\;(2.5,-0.5),\;(2,-0.5),\;$
$(1,-1),\;(0.2,-1),\;(-0.2,-1),\;(-1,-1),\;(-1.5,-0.5),\;(-2.5,-0.5),$
and
$(c,a)=$ $(1,0.2),\;(-1,0.2),\;(5,2),\;(-5,2),\;(-1,0.5)\;(1,0.5)\;(-5,0.5)\;(5,0.5),$
respectively.
@@ 7. Appendix
Table 1. The local phase portraits corresponding to Case A, which is associated with system (3.1) when $a=0,$ for Regions 1–28 of Figure 3.
@@
Table 2. The local phase portraits corresponding to Case B, which is associated with system (3.1) when $b=0,$ for Regions 1–8 of Figure 4.